A point moves in a straight line with simple harmonic motion. When it is 500mm to the right of the mean position and moving to the left, the velocity is 3m/s. When it is 1 meter to the left of the mean position and moving to the right, the velocity is 2m/s.


1
Expert's answer
2021-07-22T06:25:01-0400

"tan\phi = \frac{v(0)}{\omega x(0)} = \frac{3}{2\times0.5}"


"\phi = 71.56^0"


Amplitude :


"x_m = \frac{x(0)}{cos\phi} = \frac{1}{cos(71.56)} =3.16 m = 316 cm"